# Unsupervised Representation Learning for 3D Mesh Parameterization with Semantic and Visibility Objectives

### v1

@@ A.3 Additional Experiments and Results
Method Conformality $\uparrow$ Equiareality $\uparrow$ Ours - Semantic-Aware Param (SAMesh) 0.8694 0.6270 Ours - Semantic-Aware Param (ShDF) 0.9123 0.6707
Table 5: Ablation: effect of semantic and visibility objectives on geometry-preserving metrics. Reported are mean Conformality and Equiareality (higher is better). The numbers show that adding semantic and/or visibility objectives to the base neural UV architecture causes only minor changes in these geometry-preserving metrics while improving perceptual/qualitative behavior.
@@
Figure 13: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh) of simple meshes. We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively both partitioners produce part decompositions that align with geometric structure across simple and complex models, resulting compact, low-distortion per-part UV charts.
Figure 14: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh) of complex meshes with high number of triangle faces. We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively both partitioners produce part decompositions that align with geometric structure across simple and complex models, resulting compact, low-distortion per-part UV charts.

### v2

@@ A.3 Additional Experiments and Results
Method Conformality $\uparrow$ Equiareality $\uparrow$ Ours - Semantic-Aware Param (PartField) 0.8999 0.6921 Ours - Semantic-Aware Param (SAMesh) 0.8694 0.6270 Ours - Semantic-Aware Param (ShDF) 0.9123 0.6707
Table 7: Ablation: the effect of semantic and visibility objectives on geometry-preserving metrics. Reported are mean Conformality and Equiareality (higher is better). The Semantic-Aware and the combined Visibility+Semantic-Aware pipelines lose little equiareality relative to the baseline and even show slight improvement in conformality in some cases. Intuitively, this is because the semantic-aware stage first partitions the mesh into multiple semantically meaningful subparts and so the network can better satisfy geometry-preservation while also enforcing semantic consistency. However, the visibility-aware pipeline does not partition the mesh. Instead, it steers seams to less-visible (more occluded) regions while producing one global atlas (a single large chart). Unwrapping and optimizing a single global chart to satisfy both geometry-preservation and visibility objectives is significantly harder. Consequently, the Visibility-Aware pipeline shows a larger drop in equiareality compared to the baseline. The combined Visibility+Semantic-Aware pipeline aims to obtain the best of both worlds: (i) partition the mesh into smaller parts to retain geometry preservation, and (ii) steer seams toward less-visible regions within those parts. This combined approach yields a favorable trade-off: it loses far less equiareality (0.6369) than the pure visibility-aware pipeline (0.6093) while still improving seam visibility and semantic consistency.
@@
Table 8: Comparison of training time between our proposed semantic-aware parameterization pipeline and FlexPara’s multi-chart parameterization Zhao et al. (2025). Total times are shown outside the parentheses, and per-iteration times are shown inside.
Method Training Time (Seconds) Rabbit Potion Squidward House Average FlexPara (multi-Chart) Zhao et al. (2025) 900 (0.0900) 16377 (1.6377) 13784 (1.3784) 10353 (1.0353) Ours - Semantic-Aware Param 676 (0.1352) 666 (0.1332) 602 (0.1205) 648 (0.1296)
Table 9: Comparison of training time between our proposed visibility-aware parameterization pipeline and FlexPara’s global (single-chart) parameterization Zhao et al. (2025). Total times are shown outside the parentheses, and per-iteration times are shown inside.
Method Training Time (Seconds) Duckie Fish FlipFlop Rabbit Spot Average FlexPara (single-Chart) Zhao et al. (2025) 1306 (0.1306) 1323 (0.1323) 1275 (0.1275) 1293 (0.1293) 1230 (0.1230) 1285 (0.1285) Ours - Visibility-Aware Param 36642 (3.6642) 47460 (4.7360) 16864 (3.3729) 35232 (2.3488) 19605 (1.9605) 31160 (3.1160)
Table 10: To evaluate semantic- and visibility-awareness of the proposed method, we conducted a user‐study with 70 general participants (including graduate students with computer science and engineering backgrounds) performing 11 comparisons between textured 3D shapes and UV parameterizations produced by our method and baselines. We report the percentage of general participant preferences for each method. Our proposed method is strongly preferred by the general users over the baselines.
Method General User Preference Percentage Visibility-Awareness Evaluation OptCuts Li et al. (2018) 1.43 % FlexPara (single-Chart) Zhao et al. (2025) 7.14 % Ours - Visibility-Aware Param 91.42 % Semantic-Awareness Evaluation xatlas jpcy (2025) 4.28 % Blender Blender Foundation (2025) 4.28 % Autodesk Maya Autodesk, Inc. (2025) 10 % FlexPara (multi-Chart) Zhao et al. (2025) 7.14 % Ours - Semantic-Aware Param 74.29 %
@@
Figure 14: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh vs. PartField) of simple meshes. We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) and PartField Liu et al. (2025) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively all three partitioners can produce part decompositions that align with geometric structure across simple meshes, resulting compact, low-distortion per-part UV charts. However, it seems that PartField produces more accurate and more detailed semantic segmentations than ShDF and SAMesh in some cases (e.g., the feet and the tail are correctly detected, whereas they are not detected properly in SAMesh).
Figure 15: Ablation studies on the proposed partitioning strategy (ShDF vs. SAMesh vs. PartField) of complex meshes with high number of triangle faces. We replace our default Shape Diameter Function (ShDF) partitioner Shapira et al. (2008) with SAMesh Tang et al. (2024) and PartField Liu et al. (2025) while keeping the remainder of the partition-and-parameterize pipeline identical. Qualitatively all partitioners produce part decompositions that align with geometric structure across complex models. However, PartField appears to produce more accurate and more compact semantic segmentations than ShDF and SAMesh in some cases. For example, PartField correctly groups the entire set of fan blades into a single segmented part, whereas SAMesh and ShDF split each blade into a separate part. This is not completely wrong, but it results in too many small charts after per-part UV parameterization, which may be undesirable for some downstream tasks.
Figure 16: Examples of four of the eleven questions in our user‐study questionnaire. To evaluate semantic- and visibility-awareness of the proposed method, we conducted a user-preference study with 115 participants. We distributed the questionnaire to two groups: experts and general participants. Of the 115 participants, 70 are general participants (including graduate students with computer science and engineering backgrounds) and 45 are experts: 31 software engineers, 3 project managers, 2 product owners, 5 UV/layout artists, and 4 modeling artists, all working in the Media and Entertainment industry for film and games. Each participant completed 11 comparisons between textured 3D shapes and UV parameterizations produced by our method, FlexPara Zhao et al. (2025), OptCuts Li et al. (2018), Autodesk Maya Autodesk, Inc. (2025), Blender Blender Foundation (2025), and xatlas jpcy (2025). For each comparison, participants were rated each result according to three visual criteria: (i) texture-pattern smoothness/continuity, (ii) semantic alignment (how well colors correspond to meaningful parts), and (iii) seam visibility (how well seams are placed in occluded/less-obvious regions). The semantic-aware and visibility-aware evaluations comprised 5 and 6 questions, respectively, for a total of 11 comparisons per participant. Tables 2 and  10 report the percentage of expert and general participant preferences for each method, respectively. As the tables show, our proposed method is strongly preferred over the baselines by both expert industry-level users and general users.